The following income statement items appeared on the adjusted trial balance of Schembri Manufacturing Corporation for the year ended December 31, 2021 ($ in thousands): sales revenue, $16,300; cost of goods sold, $6,700, selling expenses. $1,350, general and administrative expenses, $850, interest revenue, $200, interest expense, $230, Income taxes have not yet been recorded. The company's income tax rate is 25% on all items of income or loss. These revenue and expense items appear in the company's income statement every year. The company's controller, however, has asked for your help in determining the appropriate treatment of the following nonrecurring transactions that also occurred during 2021 ($ in thousands). All transactions are material in amount.

 1. Investments were sold during the year at a loss of $270. Schembri also had an unrealized gain of $380 for the year on investments in debt securities that qualify as components of comprehensive income.

 2. One of the company's factories was closed during the year. Restructuring costs incurred were $1,700.

 3. During the year, Schembri completed the sale of one of its operating divisions that qualifies as a component of the entity according to GAAP. The division had incurred a loss from operations of $540 in 2021 prior to the sale, and its assets were sold at a gain of $1,500 

 4. In 2021, the company's accountant discovered that depreciation expense in 2020 for the office building was understated by $250.

 5. Negative foreign currency translation adjustment for the year totaled $300. 


Required: 

1. Prepare Schembri's single, continuous multiple-step statement of comprehensive income for 2021. Including earnings per share disclosures. One million shares of common stock were outstanding at the beginning of the year and an additional 400,000 shares were issued on July 1, 2021
Answer #1

Part 1

SCHEMBRI MANUFACTURING CORPORATION

Statement of Comprehensive Income

For the Year Ended December 31, 2021

($ in 000s)

Sales revenue

16300

Cost of goods sold

6700

Gross profit

9600

Operating expenses:

Selling expense

1350

General and administrative expenses

850

Restructuring costs

1700

Total operating expenses

3900

Operating income

5700

Other income (expense):

Loss on sales of investments

(270)

Interest expense

(230)

Interest revenue

200

Other income (expense)

(300)

Income from continuing operations before income taxes

5400

Income tax expense (25%)

1350

Income from continuing operations

4050

Discontinued operations

Income from operations of discontinued component (1500-540)

960

Income tax expense

(240)

Income from discontinued operations

720

Net income

4775

Other comprehensive income (loss)

Unrealized gains from investments, net of tax (380*(1-25%))

285

Loss from foreign currency translation, net of tax (300*(1-25%))

(225)

Total other comprehensive income

60

Comprehensive income

$4835

Earnings per share:

Income from continuing operations

$3.38

Income from discontinued operations

0.60

Net income

$3.98

Weighted average shares = 1000000+(400000/2) = 1200000

Net income from continuing operations = 4050/1200= 3.38

Net income from discontinued operations = 720/1200 = 0.60